OpenStudy (anonymous):

(x+1)(x-2)+0 solve by factoring hurry please!!!

jimthompson5910 (jim_thompson5910):

you mean (x+1)(x-2) = 0 ??

OpenStudy (anonymous):

yeah sorry

jimthompson5910 (jim_thompson5910):

set each factor equal to zero and solve

jimthompson5910 (jim_thompson5910):

if (x+1)(x-2) = 0, then x+1 = 0 which means x = ?? OR x-2 = 0 which means x = ??

OpenStudy (anonymous):

im on the phone with my teacher can you give me the answer fast

OpenStudy (dan815):

lol

OpenStudy (anonymous):

not funny like i really need it really bad

OpenStudy (dan815):

the point is to learn

OpenStudy (anonymous):

okay well thanks for the help